### v1

# Adaptive Cyclic Diffusion for Inference Scaling

Abstract
Diffusion models have demonstrated strong generative capabilities across domains ranging from image synthesis to complex reasoning tasks. However, most inference-time scaling methods rely on fixed denoising schedules, limiting their ability to allocate computation based on instance difficulty or task-specific demands adaptively. We introduce the challenge of adaptive inference-time scaling—dynamically adjusting computational effort during inference—and propose Adaptive Bi-directional Cyclic Diffusion (ABCD), a flexible, search-based inference framework. ABCD refines outputs through bi-directional diffusion cycles while adaptively controlling exploration depth and termination. It comprises three components: Cyclic Diffusion Search, Automatic Exploration-Exploitation Balancing, and Adaptive Thinking Time. Experiments show that ABCD improves performance across diverse tasks while maintaining computational efficiency.
1 Introduction
Diffusion models have become a leading class of generative models, achieving state-of-the-art performance across diverse domains, from high-fidelity image synthesis to complex language generation stable_diffusion . A core strength of diffusion models is their ability to represent complex, multimodal distributions through a hierarchical, multi-step denoising process ho2020denoising ; song2020score . This iterative refinement makes them well-suited not only for content generation but also for challenging search and reasoning tasks, such as planning controls ajay2022conditional ; chen2024diffusion , maze navigation MCTD ; zhang2025tscendtesttimescalablemctsenhanced , and puzzle solving du2022learning ; du2024learningiterativereasoningenergy .
Despite their success, a key open challenge is how to effectively realize inference-time scaling—the ability to improve model performance by allocating additional computation during inference, either on a per-task basis or, ideally, adaptively per instance. While several approaches have recently explored inference-time scaling for diffusion models li2024derivative ; ma2025inferencetimescalingdiffusionmodels ; oshima2025inferencetimetexttovideoalignmentdiffusion , they still face notable limitations.
Among the limitations, the most critical is the prevalent reliance on fixed inference procedures. Most existing methods—such as Best-of-N zhou2024diffusion , Sequential Monte Carlo singhal2025generalframeworkinferencetimescaling ; wu2023practical , Beam Search oshima2025inferencetimetexttovideoalignmentdiffusion , and Search-over-Path ma2025inferencetimescalingdiffusionmodels —execute a predetermined number of denoising steps along a trajectory. Once this trajectory completes, the inference is terminated, regardless of the complexity or specific requirements of the input instance. This rigidity fundamentally restricts the model’s ability to adapt computation based on instance difficulty, ambiguity, or the need for higher confidence or output quality. As a result, performance could suffer on more challenging inputs, and computation may be either underutilized or inefficiently spent.
In this paper, we propose a novel method—Adaptive Bi-directional Cyclic Diffusion (ABCD)—to address the challenge of adaptive inference-time scaling. This approach reframes diffusion model inference as a flexible and efficient search process, enabling adaptive computation and instance-aware refinement based on task difficulty and evolving solution quality. ABCD consists of three components. Cyclic Diffusion Search (CDS) enables iterative refinement by cycling bi-directionally through the diffusion process—alternating between denoising and re-noising steps—allowing the model to escape local minima and explore alternative generative trajectories. Automatic Exploration-Exploitation Balancing (AEEB) introduces a mechanism for adaptively controlling the depth of exploration by distributing particles across multiple re-noising levels, allowing the model to implicitly identify the appropriate amount of computation for each instance. Finally, Adaptive Thinking Time (ATT) provides a principled stopping criterion by monitoring the evolution of solution quality over inference cycles, ensuring computation is allocated efficiently and terminated when additional refinement yields diminishing returns. We demonstrate the effectiveness of ABCD across a range of tasks—including control planning, maze solving, Sudoku, and molecule generation—showing significant gains in inference-time flexibility, computational efficiency, and solution quality.
Our main contributions of this paper are as follows: (1) We formalize the problem of adaptive inference-time scaling in diffusion models and identify the limitations of fixed-step inference. (2) We propose Adaptive Bi-directional Cyclic Diffusion, a novel framework that frames inference as a flexible, search-based process with dynamic compute allocation. (3) We introduce three components—Cyclic Diffusion Search, Automatic Exploration-Exploitation Balance, and Adaptive Thinking Time—that enable iterative refinement, adaptive search control, and automatic stopping. (4) We empirically validate ABCD across diverse tasks, including planning, maze solving, Sudoku, and molecule generation, showing improvements in flexibility, accuracy, and compute efficiency.
2 Preliminaries
Diffusion model. Diffusion models ho2020denoising ; sohl2015deep ; song2022denoisingdiffusionimplicitmodels ; song2020score have recently emerged as a powerful generative framework that formulates sample generation as a gradual denoising process. This framework consists of two complementary components: a fixed forward process, which incrementally injects noise into data to transform structured inputs into nearly pure noise, and a learned reverse process, which aims to recover clean data from noisy observations. Formally, the forward process defines a sequence of latent variables $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ conditioned on the original sample $\mathbf{x}_{0}$ via a Markov chain: $q(\mathbf{x}_{t+1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t+1};\sqrt{\alpha_{t}}\mathbf{x}_{t},(1-\alpha_{t})\mathbf{I}),$ where $\alpha_{t}\in(0,1)$ denotes a pre-defined noise schedule. This process ultimately transforms $\mathbf{x}_{0}$ into Gaussian noise $x_{T}\sim\mathcal{N}(0,I)$. The forward transition at an arbitrary step $t$ is given by:
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},\quad\mathbf{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),$
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ denotes the cumulative product of the noise schedule. This expression provides the marginal distribution of $\mathbf{x}_{t}$ conditioned on $\mathbf{x}_{0}$.
The reverse process uses a neural network to predict the noise component $\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},t)$, which is then used to estimate the clean sample: $\mathbf{\hat{x}}_{0}(\mathbf{x}_{t})=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right).$ Using this estimate, the previous latent state is computed via update:
$\mathbf{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\,\mathbf{\hat{x}}_{0}(\mathbf{x}_{t})+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},t)+\sigma_{t}\mathbf{z},\quad\mathbf{z}\sim\mathcal{N}(0,\mathbf{I}),$
where $\sigma_{t}$ controls the level of stochasticity. The reverse process can be parameterized in various equivalent forms, including noise prediction $\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},t)$, posterior mean prediction $\mathbf{\mu}_{\theta}(\mathbf{x}_{t},t)$, or direct clean-sample prediction $\mathbf{\hat{x}}_{0}(\mathbf{x}_{t})$.
Reward-guided generation. In many practical scenarios, it is desirable not only to generate realistic samples from a diffusion model, but also to align them with a predefined reward function $r(\mathbf{x})\in\mathbb{R}$. A common objective is to sample from a distribution that biases toward high-reward samples while staying close to the original model distribution uehara2025inferencetimealignmentdiffusionmodels . This can be formalized as:
$p^{(\alpha)}(\mathbf{x})\propto\exp(r(\mathbf{x})/\alpha)\cdot p_{\text{pre}}(\mathbf{x})$,
where $\alpha>0$ is a temperature parameter controlling the trade-off between reward optimization and fidelity to the pre-trained distribution $p_{\text{pre}}$. As $\alpha\to 0$, the distribution becomes increasingly concentrated on reward-maximizing samples li2024derivative .
Figure 1: Overview of our method. (1) Start with $N$ particles, use jumpy denoising $x_{T}\rightarrow x_{T-j}\rightarrow\cdot\cdot\cdot\rightarrow x_{0}$. (2) Replicate each particles $J$ times and send each particles to multiple noise levels (AEEB). (3) Re-denoise particles. (4) Select Top $K$ particles. (5) We repeat steps (2)–(4) until the adaptive terminal condition is satisfied (CDS). The number of cycles executed in this way defines the Adaptive Thinking Time (ATT).
3 ABCD: Adaptive Bi-directional Cyclic Diffusion
We propose Adaptive Bi-directional Cyclic Diffusion (ABCD), a novel inference framework for diffusion models that enables adaptive and efficient scaling with respect to the available computation budget. ABCD is designed to flexibly trade off exploration and exploitation during inference and to dynamically determine when to terminate. The framework is composed of three core components: (i) Cyclic Diffusion Search, (ii) Automatic Exploration, and (iii) Adaptive Thinking Time.
3.1 Cyclic Diffusion Search
The core idea of Cyclic Diffusion Search is to enable iterative inference by cycling bi-directionally through the diffusion timeline, alternating between the denoising (reverse) and noising (forward) processes. This approach allows the model to refine its predictions over time while adaptively revisiting earlier stages of the generative process to explore alternative solution paths. Each iteration of the cycle consists of three main stages: Denoising, Selection-and-Copy, and Noising.
Denoising. As a particle-based method, the inference process begins by sampling a set of $N$ particles from the standard Gaussian prior distribution ${\cal N}(0,{\bf I})$. Leveraging the pretrained diffusion model, each particle is then denoised from $t=T$ to $t=0$ using a coarse, accelerated Denoising Diffusion Implicit Models (DDIM) (song2022denoisingdiffusionimplicitmodels, ) trajectory that skips intermediate steps. This allows us to quickly obtain initial sample estimates ${\bf x}_{0}$ from the noisy latents ${\bf x}_{T}$, significantly reducing the time to acquire a preliminary solution candidate. While this fast DDIM process may yield suboptimal initial outputs, the key idea of our framework is to mitigate this through subsequent cyclic refinement. Unlike some prior approaches (ma2025inferencetimescalingdiffusionmodels, ) that interleave intermediate noising steps during the denoising pass, we prioritize reaching $t=0$ quickly to establish a diverse set of starting points for the cyclic search.
Selection and Copy. Once all particles reach $t=0$, we evaluate them using a task-specific reward function, yielding scores ${r_{1},r_{2},\dots,r_{N}}$. To guide the subsequent search towards promising regions, we select the top-$K$ particles with the highest reward scores. These $K$ selected particles serve as “anchor” points for the next phase of exploration. Each of these $K$ particles is then replicated $J$ times, producing a total of $K\times J$ particles.
Noising. The replicated particles are then sent back to an earlier diffusion step $t^{\prime}>0$—referred to as the go-back timestep—via the noising process of the diffusion model: $q({\bf x}_{t^{\prime}}|{\bf x}_{0})$. This operation reintroduces stochasticity and enables the particle to re-enter the denoising trajectory from a different point in the diffusion space. The full cycle—denoising, selection-copy, and noising—is then repeated iteratively until a predefined stopping criterion is met.
3.2 Automatic Exploration-Exploitation Balancing
A key limitation of the basic Cyclic Diffusion Search, as described above, lies in its reliance on a fixed go-back timestep, treated as a manually specified hyperparameter. In principle, there may exist a go-back step that optimally balances exploration and exploitation at each stage of inference. However, identifying this optimal step is generally difficult and time-consuming. Moreover, our experiments show that the optimal go-back step not only varies across different instances but could also evolve dynamically throughout the inference trajectory of a single instance. Consequently, relying on a fixed go-back step is both hard to tune and fundamentally suboptimal for adaptive inference.
Illustrative example. Consider solving a Sudoku puzzle. In the early stages, when only a few digits are provided, global exploration is beneficial to fill in plausible candidates and explore a wide solution space. In contrast, in the later stages—when most of the board is correctly filled—aggressive global changes can corrupt nearly-complete solutions. At that point, cautious local refinement is preferable. This illustrates the need to dynamically adjust the exploration-exploitation balance during inference.
Implementation. To implement automatic exploration mechanism, we modify the noising step of the previously described cyclic diffusion search as follows. We begin by defining a set of predefined go-back timesteps, referred to as the “temperature pool”, denoted by $\mathcal{T}=(t_{1},t_{2},\dots,t_{M})$, where each $t_{i}$ corresponds to a distinct level of exploration. A simple and effective way to construct $\mathcal{T}$ is to uniformly partition the diffusion timeline into $M$ segments. After the select-and-copy step, yielding $K\times J$ particles, the replicas are distributed across the predefined temperatures in $\mathcal{T}$ via the forward (noising) process, rather than being sent to a single go-back timestep. This allows each anchor particle to probe multiple exploration depths in parallel, enabling the system to implicitly identify effective go-back steps through iterative feedback. After completing the re-noising, all $K\times J$ particles are denoised back to $t=0$, and a new top-$K$ set is selected. This way, particles from suboptimal temperatures will automatically be discarded. This cycle continues until a stopping criterion is satisfied. See Figure 1 for an illustration of the algorithm.
Algorithm 1 ABCD: Adaptive Bi-directional Cyclic Diffusion
1:procedure ABCD(verifier, $T$, $\mathcal{T}$, max_iter, $\kappa$, $N$, $K$, $J$)
2:     Initialize $N$ particles $\{x_{T}^{(i)}\}$ from ${\cal N}(0,I)$
3:     Denoise to obtain $\{x_{0}^{(i)}\}$ via Fast denoising
4:     while less than max_iter do
5:         Selection: Select top-$K$ particles using verifier
6:         If Selected top-$K$ particles come from smallest $t_{g}$ for $\kappa$ times consecutively
7:          break
8:         Copy: Replicate each $K$ particles $J$ time
9:         Noising: Send each $K$ particles to each $t_{g}\in$ $\mathcal{T}$
10:         Denoising: Denoise from each $t_{g}$ go-back temperature
11:     end while
12:     return Select best particle from top-$K$ particles
13:end procedure
3.3 Adaptive Thinking Time
When should we stop this cyclic inference computation? Unlike single-pass inference methods that terminate once reaching $t=0$, ABCD’s cyclic structure raises a nontrivial question of when to conclude the iterative search. A naive solution is to impose a fixed computation budget (e.g., a maximum number of cycles), but this risks either underutilizing available compute or wasting resources when further refinement yields no benefit.
To address this, ABCD incorporates an implicit uncertainty measure derived from the temperature distribution of the top-$K$ particles. After each cycle, we record the re-noising temperatures from which the current top-$K$ particles originated. If high-temperature particles are present in the top-$K$, this suggests that exploration is still yielding valuable improvements. Conversely, if all top-$K$ particles originate from the lowest temperature, it implies that global exploration has little impact and the search is now confined to local refinement—signaling that the process may have converged.
As a practical implementation of this idea, we adopt a simple yet effective stopping rule: terminate the inference when all top-$K$ particles originate from the lowest temperature for $\kappa$ consecutive cycles. The hyperparameter $\kappa$ controls the sensitivity of the termination criterion, balancing early stopping against the opportunity for further refinement.
4 Related Works
Inference-Time Scaling with Pre-Trained Diffusion Models. There is growing interest in leveraging pre-trained diffusion models more effectively during inference, by adaptively steering the generation process during sampling. A unified perspective brekelmans2025posteriorinference ; uehara2025inferencetimealignmentdiffusionmodels frames this as sampling from a reward-augmented distribution of the form $p(\mathbf{x})\exp(\lambda r(\mathbf{x}))$. A common approach is classifier guidance dhariwal2021diffusion ; ho2022classifierfreediffusionguidance , which assumes access to a differentiable value function. For non-differentiable rewards, derivative-free strategies such as SVDD li2024derivative , Beam Search oshima2025inferencetimetexttovideoalignmentdiffusion , DSearch li2025dynamicsearchinferencetimealignment , and MCTS zhang2025tscendtesttimescalablemctsenhanced have been proposed. Alternatively, sampling-based approaches like Importance Sampling and Sequential Monte Carlo (e.g., TDS wu2023practical , FKD singhal2025generalframeworkinferencetimescaling ) offer principled reward-weighted trajectory selection. However, most methods operate under uni-directional denoising and rely on static search schedules.
Planning and Reasoning with Diffusion Models. Diffusion models have recently gained traction for long-horizon planning and complex reasoning tasks. Diffuser janner2022planning reinterprets planning as conditional generation over offline trajectories. Subsequent works ajay2022conditional ; chen2024diffusion ; ding2024diffusionworldmodelfuture ; zhou2024diffusion extend this idea by using diffusion models as world models. On the reasoning side, recent studies explore diffusion models for structured inference. Energy-based diffusion models have been applied to symbolic reasoning tasks du2022learning ; du2024learningiterativereasoningenergy ; zhang2025tscendtesttimescalablemctsenhanced , and SRM wewer2025spatial applies diffusion to visual spatial reasoning.
5 Experiments
We evaluate our proposed method on a diverse suite of challenging tasks that require generating samples in sparse or previously unexplored regions at training phase: (1) toy Mixture of Gaussian(MoG), (2) point-mass maze navigation (park2024ogbench, ), (3) Sudoku puzzle completion (palm2018recurrentrelationalnetworks, ; wang2019satnetbridgingdeeplearning, ), (4) path generation on unseen pixel maze image (maze, ), and (5) molecular structure prediction (hoogeboom2022, ). These tasks present distinct challenges: MoG shows the importance of multiple go-back temperature pool as the proof-of-concept; point-mass maze requires long-horizon planning over 1000 steps; Sudoku demands logical consistency across row, column, and block constraints; path generation on unseen maze image tests generalization to novel environmental structures; and molecular structure prediction requires generating valid 3D conformations under chemical and physical constraints.
5.1 Baselines
We compare ABCD against several strong representative baselines: Base Diffusion: The standard diffusion model without additional computational scaling, serving as our primary baseline. Best-of-$N$ (BoN): A computationally scaled variant that generates $N$ independent samples and selects the best according to a task-specific verifier used consistently across all models. Diffusion Beam Search (BS) (oshima2025inferencetimetexttovideoalignmentdiffusion, ): Combines diffusion with beam search to scale the denoising process, pruning unpromising candidates during generation. Sequential Monte Carlo Diffusion (SMC) (singhal2025generalframeworkinferencetimescaling, ; wu2023practical, ): Employs resampling during the denoising process to focus computational resources on more promising trajectories. Implementation details were adapted from  singhal2025generalframeworkinferencetimescaling to suit our experimental setting. Search over Paths (SoP) (ma2025inferencetimescalingdiffusionmodels, ): Enhances generation by allowing limited backward steps (adding noise) to promising samples, enabling exploration of wider solution spaces compared to strictly unidirectional approaches.
All baselines used the same pre-trained diffusion model per tasks, and inference-time budgets were expanded as much as possible while preserving their core operational characteristics. In BS and SoP, inference-time compute was increased by more frequent expansion and selection, while SMC scaled compute via higher resampling frequency, and BoN by increasing the number of particles $N$. For fair comparison, the total number of particles was set equal across all methods (except BoN).
Figure 2: Evaluated Tasks. We evaluate ABCD on one toy tasks and four challenging tasks requiring generation in sparse regions of solution space: (a) Mixture of Gaussian for Proof-of-concept that we need multiple go back temperature. (b) Sudoku puzzle completion demanding logical consistency across multiple constraints simultaneously; (c) Pixel Maze Path Finding testing generalization to novel environmental structures not encountered during training; (d) Molecular structure prediction requiring physically and chemically valid 3D conformations. (e) OGBench PointMaze navigation requiring complex planning over 1000+ steps;
5.2 Mixture of Gaussian
Result. We evaluate ABCD against four fixed single go-back temperatures (GB 20, 40, 60, 80) under two inference regimes: a fixed 25-cycle budget and an open-ended budget of up to 100 cycles with early stopping upon goal attainment. As shown in comparison with fixed 25 cycles (Table 2): only ABCD attains 100 % success and minimal final distance on both datasets. Fixed small go-back steps work on Dataset 1 but cannot escape poor initializations in Dataset 2, whereas fixed large steps succeed escaping from poor initializations on Dataset 2 but over-cycle on Dataset 1 and 2 (e.g., reaching the goal at cycle 20 but diverging by cycle 25 after losing useful information). As shown in comparison with up to 100 cycles with early stopping (Table  2): some fixed levels solve one dataset, but only ABCD consistently achieves both 100 % success with comparable average cycles. These findings underscore that exploration (coarse jumps) and exploitation (fine steps) require different go-back levels—a balance that ABCD’s adaptive termination naturally realizes.
Table 1: Left: Success rate (%) and distance to the goal on MoG. We compare multi go-back (ABCD) and fixed single go-back (GB 20, 40, 60, 80).
Method Dataset 1 Dataset 2
Success Distance Success Distance
ABCD (Ours) 100 0.08 ± 0.0 100 0.03 ± 0.0
GB 20 95 0.17 ± 0.0 26 3.97 ± 6.0
GB 40 0 2.90 ± 0.6 59 1.73 ± 4.6
GB 60 0 7.80 ± 0.8 7 2.94 ± 1.1
GB 80 0 42.81 ± 6.6 0 36.46 ± 11.1
Table 2: Right: Success rate (%) and mean iteration to reach the goal on MoG. Every method terminates when it reaches the goal mode.
Method Dataset 1 Dataset 2
Success Mean Cyc Success Mean Cyc
ABCD (Ours) 100 3.36 100 7.02
GB 20 100 2.06 79 16.14
GB 40 8 46.38 100 6.77
GB 60 0 100 15.05
GB 80 0 6 43.33
5.3 Sudoku puzzle completion
Figure 3: Mean accuracy across four difficulty levels—Very Hard (17-19 provided), Hard (20-21 provided), Medium (22-24 provided), and Easy (25-27 provided).
Result. Figure 4(Left) shows that ABCD consistently outperforms all baselines, achieving higher accuracy in less wall-clock time and displaying superior inference-time scaling. Notably, SoP (ma2025inferencetimescalingdiffusionmodels, ) never achieves perfect accuracy (95.5 % accuracy), whereas ABCD attains 100 % accuracy in fewer cycles. Moreover, as illustrated in Figure 3, ABCD’s accuracy remains stable across all difficulty levels, while baseline methods exhibit marked performance degradation on the hardest instances. We attribute this robustness to ABCD’s dynamic computation allocation: it adaptively concentrates additional inference effort on more challenging puzzles, in contrast to baselines that expend a fixed, uniform budget.
Figure 4: Left: Sudoku Puzzle Completion result. Mean accuracy on Harder dataset (17-28 entities provided) by giving more computational budget (sec.). Middle: Pixel maze path finding result. Success rate on the OOD Pixel Maze size-15 test set. Right: Molecular 3D structure prediction task result. Molecular Stability rate on the QM9 dataset. For each method, inference-time computation was scaled as much as possible along its core controllable axis for inference time scaling to ensure a fair comparison under expanded budgets.
5.4 Path generation on unseen pixel maze image
Setup. We evaluated ABCD’s adaptability on a path-generation task using mazes of previously unseen sizes (maze, ). Every method leveraged the same diffusion model pretrained on small mazes (sizes 4, 5, and 6) and was then tested on larger, structurally distinct mazes zhang2025tscendtesttimescalablemctsenhanced . This setting rigorously probes each algorithm’s capacity to generalize to novel spatial configurations at inference time.
Figure 5: Pixel Maze path finding results on multiple maze size. Success rate comparison across maze sizes in the Pixel Maze task. All settings involve OOD evaluation, where models are trained on smaller mazes (sizes 4$\sim$6) and tested on larger ones. As maze size increases, the search space grows and the performance gap between other methods widens, highlighting the importance of adaptive inference-time exploration in complex scenarios.
Result. As shown in Figure 4(Middle), ABCD reaches near-perfect success rates significantly faster than all baselines, evidencing a superior time–accuracy trade-off. In contrast, most competing methods—including Base Diffusion, BoN, BS, and SMC—never exceed a 60 % success rate even under generous computational budgets. ABCD, by comparison, attains 100 % success in substantially less time. Figure 5 further examines performance across maze sizes 11–15 in the Pixel Maze task. Here again, ABCD consistently dominates both success rate and inference time at every size, and the margin over baselines widens as the search space grows. These findings underscore the value of adaptive, instance-specific exploration strategies especially in complex search demanding tasks.
5.5 Molecule generation
Setup. We evaluate ABCD on the 3D molecule generation task using the QM9 dataset ramakrishnan2014quantum , which contains 130k small molecules with annotated properties and atomic coordinates. For our evaluation, we utilized the Equivariant Diffusion Model (EDM) hoogeboom2022 as the pre-trained model. We then applied various inference-time methods to generate valid 3D conformations.
Result. As shown in the Figure 4(Right), ABCD significantly outperforms all baselines by achieving higher molecular stability with less inference wall-clock time. For instance, ABCD reaches a molecule stability of 0.97 within approximately 20 seconds, a threshold that other baselines either fail to attain or require considerably more computational time to approach. Notably, our method achieved a peak molecule stability of approximately 0.99, which is about 5.32% higher than the peak stability of 0.94 achieved by SoP (ma2025inferencetimescalingdiffusionmodels, ), the next best-performing baseline. Furthermore, the performance gap between ABCD and the baselines widens at higher stability levels, emphasizing the robustness of our method in complex molecular generation tasks.
Figure 6: OGBench PointMaze task results. Success rates across large and giant mazes. Our ABCD approach consistently outperforms baselines, rapidly achieving higher success rates with fewer inference time and attaining perfect performance, particularly in the Giant maze, emphasizing its superior efficiency.
5.6 OGBench Point Maze
Setup. We further evaluate the ABCD framework using OGBench’s PointMaze environments park2024ogbench . In this task, agents must navigate to a specified goal region. The base diffusion model follows Diffuser janner2022planning , and is trained using the offline dataset provided by OGBench park2024ogbench , containing trajectories collected by a noisy expert policy. Each environment includes 5 distinct tasks with different start-goal pairs. At inference time, the goal information is provided to the verifier for scoring the particles. To let the model efficiently plan over long horizon while reaching the goal as quickly as possible, we employed a value guidance function dhariwal2021diffusion to minimize the distance from the final state to the goal.
Result. As shown in Figure 6, ABCD consistently surpasses other baselines. Moreover, only ABCD achieves perfect performance on both mazes. It reflects the advantages of our method: by quickly denoising to $t=0$, it locates valid trajectories almost immediately, rather than spending compute on uniformly long reverse sampling path. In maze giant, the gap widens because larger mazes demand more on both global exploration (to reach the goal) and fine refinement (to thread tight corridors).
5.7 Do we need to care about go-back temperature level?
Figure 7: (a-b) Per-instance analysis of inference-time scaling behavior across different go-back temperatures. (a) Results for Sudoku instances with 14 provided entities. (b) Results for instances with 44 provided entities. (c) Performance comparison with vs. without adaptive terminal condition in Pixel Maze Size 15. x-axis is the average inference time per sample. (d) Different thinking assignment per hardness in Sudoku (provided number from 17 to 34).
Instance-Level optimal temperature. Figure 7(a–b) plots cycle-wise accuracy for two Sudoku instances (14 vs. 44 givens) under fixed go-back noise levels. Certain noise settings plateau after only a few cycles and the plateau point differs between puzzles, indicating that each instance possesses its own optimal noise level.
Temperature selection per cycle. Figure 8 tracks the provenance of selected particles source across cycles for three instance types in Sudoku (17, 19, and 32 givens). Harder puzzles continue sampling from a broad range of noise levels, while easier ones converge rapidly to low-variance perturbations. Early iterations emphasize high-variance (global) moves; later iterations favor low-variance (local) refinements. These patterns demonstrate ABCD’s dynamic, per-instance generative process through adaptive exploration vs. exploitation.
Figure 8: The origin of the top-10 selected particles at each cycle. The above plot shows, for different Sudoku difficulty cases, the origin of the Top-10 selected particles at each cycle. We observe that the behavior varies significantly across cases, and harder instances tend to require longer search. Additionally, early iterations tend to select samples with large modifications, while later cycles focus on smaller refinements. This indicates that our model automatically expands and adjusts the diffusion generation process, resulting in a different generation graph for each final $\textbf{x}_{0}$ prediction.
5.8 Do we need adaptive terminal condition?
Efficiency of adaptive terminal condition. An ablation study on Pixel maze size 15 (Figure 7(c) compares fixed cycle length inference (blue) to our adaptive terminal criterion (red) with the same temperature pool. The adaptive policy stops early on simple puzzles—avoiding wasted compute—yet allocates extra iterations to hard ones, yielding consistently superior time–success trade-offs.
ABCD automatically assigns proper thinking time. Figure 7(d) demonstrates that ABCD allocates more thinking cycles to harder puzzles (fewer givens) and terminates earlier on simpler ones, validating that our adaptive terminal criterion scales compute with problem difficulty. Furthermore, the stop-flag hyperparameter $\kappa$ directly controls exploration depth: larger $\kappa$ values enforce additional cycles before stopping, which boosts success on challenging instances. Identical trends hold on the Pixel Maze task.
6 Limitations and Discussion
While ABCD performs well on moderately sized planning and reasoning tasks, its scalability to high-dimensional output spaces—such as high-resolution image or video generation—remains unexplored. Extending ABCD to such domains may require structural priors or hierarchical variants to maintain tractability and efficiency. In addition, although ABCD demonstrates strong empirical performance through its search-based inference strategy, its theoretical properties remain largely unexamined. A deeper theoretical understanding could be gained by connecting ABCD to principled frameworks such as Bayesian optimization or approximate inference.
A particularly promising future direction lies in amortizing the search process. The optimal re-noising steps and decisions discovered during ABCD’s iterative inference could be treated as pseudo-labels for training a context-aware policy that predicts adaptive search strategies conditioned on the input. Such a learned policy could retain the adaptivity of ABCD while significantly reducing inference-time cost by eliminating online search. We leave the development of such amortized inference mechanisms to future work.
7 Conclusion
We introduced Adaptive Bi-directional Cyclic Diffusion (ABCD), a novel inference framework that enables instance-aware, compute-adaptive generation with diffusion models. By framing inference as a flexible search process, ABCD dynamically allocates computation through bi-directional diffusion cycles, adaptive exploration, and principled stopping. Across diverse domains—including planning, reasoning, and molecule generation—ABCD improves performance and efficiency over fixed-schedule baselines. This work underscores the promise of adaptive inference-time scaling.
Acknowledgments and Disclosure of Funding
We thank to Junyeob Baek, Doojin Baek, and Hyeonseo Cho for insightful discussions and assistance with this project. This research was supported by GRDC (Global Research Development Center) Cooperative Hub Program (RS-2024-00436165) and Brain Pool Plus Program (No. 2021H1D3A2A03103645) through the National Research Foundation of Korea (NRF) funded by the Ministry of Science and ICT. The authors of Mila acknowledge the support from CIFAR, NSERC, the Future of Life Institute.